Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
gherardini - 6:56 Tuesday 28 October 2025 (68046) Print this report
Operator Report - Night shift
Even restoring of PR CHRoCC DAC value, done in the afternoon, it didn't go back to its initial condition in term of temperature, voltage and current values; during the night I tried a couple of time to go on with the lock acquisition with the SR misaligned but the ITF failed in the same way of yesterday night, unlocking while started to lock the OMC because of too much power on the dark fringe, I left the ITF locked at CARM_NULL_1F for the night.
Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×